# Directional Sheaf Hypergraph Networks: Unifying Learning on Directed and Undirected Hypergraphs

### v1

@@ A.1 Spectral Properties
By definition, we have $L^{\vec{\mathcal{F}}}_{N}:=I_{nd}-Q^{\vec{\mathcal{F}}}_{N}$, with $Q^{\vec{\mathcal{F}}}_{N}:=D_{V}^{-\frac{1}{2}}{B^{(q)}}^{\dagger}D_{E}^{-1}B^{(q)}D_{V}^{-\frac{1}{2}}$.
$Q^{\vec{\mathcal{F}}}_{N}$ can be factored as
$Q^{\vec{\mathcal{F}}}_{N}=\left(D_{V}^{-\frac{1}{2}}{B^{(q)}}^{\dagger}D_{E}^{-\frac{1}{2}}\right)\left(D_{E}^{-\frac{1}{2}}B^{(q)}D_{V}^{-\frac{1}{2}}\right)=\left(D_{E}^{-\frac{1}{2}}B^{(q)}D_{V}^{-\frac{1}{2}}\right)^{\dagger}\left(D_{E}^{-\frac{1}{2}}B^{(q)}D_{V}^{-\frac{1}{2}}\right).$
@@
Since $L^{\vec{\mathcal{F}}}_{N}:=I_{nd}-Q^{\vec{\mathcal{F}}}_{N}$, it follows that the spectrum of $L^{\vec{\mathcal{F}}}_{N}$ is upper-bounded by 1, which concludes the proof. ∎
@@ A.2 Generalization Properties
In the $2$-uniform case, every hyperedge $e$ contains exactly two nodes (i.e., $\delta_{e}=2$). Consider the general expression of the unnormalized Laplacian given in Eq. 3. Since the graph has no directions, $\mathcal{S}^{(0)}_{u\trianglelefteq e}=1$ for all $u\in V,e\in E$, and for any choice of the charge parameter $q$. As a result, the off-diagonal terms of $L^{\vec{\mathcal{F}}}$ are real-valued (the diagonal ones always are).
In particular, when $\delta_{e}=2$ for all $e\in E$, $L^{\vec{\mathcal{F}}}$ reads:
$(L^{\vec{\mathcal{F}}})_{uv}=\begin{cases}\displaystyle\tfrac{1}{2}\sum_{e:\,u\in e}\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e}\in\mathbb{R}^{d\times d},&u=v,\\[10.0pt] \displaystyle-\tfrac{1}{2}\sum_{e:\,u,v\in e}\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\in\mathbb{R}^{d\times d},&u\neq v.\end{cases}$
Thus, $L^{\vec{\mathcal{F}}}$ precisely coincides with the Sheaf Laplacian of Hansen and Gebhart [2020] up to the multiplicative constant $\tfrac{1}{2}$.
When considering the case of a trivial Sheaf (i.e., when $\mathcal{F}_{v\trianglelefteq e}=1$), $L^{\vec{\mathcal{F}}}$ coincides with the definition of the classical graph Laplacian L=DAL=D-A, where $A$ is the adjacency matrix and $D$ is the node degree matrix.
@@
The Magnetic Laplacian proposed by Zhang et al. [2021] is defined as
$L^{(q)}\coloneqq D_{s}-H^{(q)}=D_{s}-A_{s}\odot\exp\!\bigl(i\,\Theta^{(q)}\bigr),$
where $\Theta^{(q)}$ denotes the phase matrix defined as
$\Theta^{(q)}\coloneqq\exp\bigl(2\pi q\,(A-A^{\top})\bigr)$
and $A_{s}$ is the symmetrized adjacency matrix defined as
$A_{s}:=\tfrac{1}{2}(A+A^{\top})$
and $D_{s}$ is a diagonal matrix defined as
$(D_{s})_{uu}:=\sum_{v\in V}(A_{s})_{uv}\text{ for all $u\in V$}.$
Entry-wise, $H^{(q)}$ can be written as:
$H^{(q)}_{uv}=\begin{cases}\tfrac{1}{2}e^{\,2\pi iq}&(u,v)\in E\\[3.0pt] \tfrac{1}{2}e^{-\,2\pi iq}&(v,u)\in E\\[3.0pt] 1&\{u,v\}\in E\\[3.0pt] 0&\text{otherwise}.\end{cases}$
In the directed, 2-uniform case, every hyperedge $e$ contains exactly two nodes ($\delta_{e}=2$). For every $e\in E$, the product $(\mathcal{S}^{(q)}_{u\trianglelefteq e})^{\dagger}\mathcal{S}^{(q)}_{v\trianglelefteq e}$ can take one of the following three values:
@@
    $\mathcal{S}^{(q)}_{u\trianglelefteq e}=\mathcal{S}^{(q)}_{v\trianglelefteq e}=1\quad\Longrightarrow\quad\big(\mathcal{S}^{(q)}_{u\trianglelefteq e}\big)^{\dagger}\mathcal{S}^{(q)}_{v\trianglelefteq e}=1.$
@@
    $\mathcal{S}^{(q)}_{u\trianglelefteq e}=e^{-2\pi iq},\qquad\mathcal{S}^{(q)}_{v\trianglelefteq e}=1\quad\Longrightarrow\quad\big(\mathcal{S}^{(q)}_{u\trianglelefteq e}\big)^{\dagger}\mathcal{S}^{(q)}_{v\trianglelefteq e}=e^{+2\pi iq}.$
@@
    $\mathcal{S}^{(q)}_{u\trianglelefteq e}=1,\qquad\mathcal{S}^{(q)}_{v\trianglelefteq e}=e^{-2\pi iq}\quad\Longrightarrow\quad\big(\mathcal{S}^{(q)}_{u\trianglelefteq e}\big)^{\dagger}\mathcal{S}^{(q)}_{v\trianglelefteq e}=e^{-2\pi iq}.$
@@
$(Q^{\vec{\mathcal{F}}})_{uv}=\begin{cases}\begin{aligned} \frac{1}{2}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\,\vec{\mathcal{F}}_{v\trianglelefteq e}&=\frac{1}{2}\,e^{+\,2\pi iq}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}=\frac{1}{2}e^{+\,2\pi iq},\end{aligned}&\text{if }e=(u,v),\\[12.0pt] \begin{aligned} \frac{1}{2}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\,\vec{\mathcal{F}}_{v\trianglelefteq e}&=\frac{1}{2}\,e^{-\,2\pi iq}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}=\frac{1}{2}\,e^{-\,2\pi iq},\end{aligned}&\text{if }e=(v,u),\\[12.0pt] \begin{aligned} \frac{1}{2}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\,\vec{\mathcal{F}}_{v\trianglelefteq e}&=\frac{1}{2}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}=1,\end{aligned}&\text{if }e=\{u,v\}.\end{cases}$
@@
$Q^{\vec{\mathcal{F}}}={B^{(q)}}^{\dagger}\,D_{E}^{-1}\,B^{(q)}\;=\;H^{(q)},\text{ with }D_{V}\;=\;D_{s}.$
@@
$L^{\vec{\mathcal{F}}}=D_{s}-H^{(q)}=L^{(q)}.$
Lastly, noticing that, by construction, the Sign-Magnetic Laplacian proposed in Fiorini et al. [2023] coincides with the Magnetic Laplacian when $q=\frac{1}{4}$, we conclude that our operator also generalizes the former. ∎
@@
In the unit-weight case, the Laplacian matrix proposed by Zhou et al. [2006] for undirected hypergraphs is defined as follows:
$\Delta:=I-Q_{N}\qquad\text{ with }Q_{N}:=D_{V}^{-\frac{1}{2}}BD_{E}^{-1}B^{\top}D_{V}^{-\frac{1}{2}}.$
Since any undirected hypergraph be regarded as a special case of a directed hypergraph in which every hyperedge consists solely of tail nodes (or, equivalently, solely of head nodes), as shown in Eq. 4, in our proposed Laplacian matrix $L^{\vec{\mathcal{F}}}_{N}$ each product of two restriction maps reduces to a real weight of $1$, therefore contributing only to the real part of the operator. In particular, for a trivial sheaf where $\mathcal{F}_{v\trianglelefteq e}=1$, the incidence matrix $B^{(q)}$ in Eq. 1 reduces to the transpose of the binary incidence matrix $B$ of Zhou et al. [2006]. ∎
@@
Our Normalized Directed Sheaf Hypergraph Laplacian $L^{\vec{\mathcal{F}}}_{N}$, component-wise reads:
$(L^{\vec{\mathcal{F}}}_{N})_{uv}=\left\{\begin{array}[]{lr}\displaystyle\par I_{d}-D_{u}^{-1}\sum_{e:u\in e}\frac{1}{\delta_{e}}{\mathcal{F}}_{u\trianglelefteq e}^{\top}{\mathcal{F}}_{u\trianglelefteq e}&u=v\par\\ \displaystyle-D_{u}^{-\frac{1}{2}}\big(\sum_{e:u,v\in e}\frac{1}{\delta_{e}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\big)D_{v}^{-\frac{1}{2}}&u\neq v.\end{array}\right$.
@@
$(L^{\vec{\mathcal{F}}}_{N})_{uv}=\left\{\begin{array}[]{ll}\displaystyle 1-\sum_{e:\,u\in e}\frac{1}{d_{u}\,\delta_{e}},&u=v,\\[6.0pt] \displaystyle-\!\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u,v\in H(e)\\ \lor\;u,v\in T(e)\end{subarray}}\!\!\!\!\!\frac{1}{\delta_{e}}-\mathbf{i}\left(\sum_{\begin{subarray}{c}e\in E\\ u\in T(e)\\ \land v\in H(e)\end{subarray}}\!\!\frac{1}{\delta_{e}}-\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u\in H(e)\\ \land v\in T(e)\end{subarray}}\!\!\frac{1}{\delta_{e}}\right)\frac{1}{\sqrt{d_{u}}\sqrt{d_{v}}},&u\neq v.\end{array}\right$. (13)
Such an expression coincides with the definition of the Generalized Directed Laplacian when considering $W=I$. ∎
@@ B.1 Impact of charge parameter
The charge parameter $q$ controls how much each directed hyperedge contributes to the real and imaginary parts of the Directed Sheaf Hypergraph Laplacian. Larger values of $q$ place more directional information in the imaginary component, whereas smaller values reduce the directional contribution, emphasizing orientation-agnostic interactions in the real part. Because the dataset differ in how informative directionality is, the optimal $q$ is inherently data-dependent. In practice, a careful tuning of it is needed $q$ to select the value that yields the best performance, allowing either a partial or a full contribution of directional information to be encoded as needed. Table 2 reports the values chosen by our hyperparameter tuning procedure. As one can seen, for most datasets the hyperparameter tuning procedure sets a relatively high importance to directional information for each dataset, particularly for Telegram and Roman-empire. On the other hand, as observed in Zhang et al. [2021], in highly homophilic networks such as Citeseer and Cora the direction of hyperedges impacts negatively on the performance. Synthetic datasets instead interestingly seem to see a better performance with intermediate values of $q$, which could further explain the difference in performance with other directed methods such as GeDi-HNN and DHGNN.
@@ B.2 Impact of stalk dimension and number of layers
As noted in Section 1, standard HGNNs are prone to oversmoothing: as network depth increases, node representations become indistinguishable and accuracy degrades. In Fig. 4, we study how depth and the stalk dimension $d$ affect the accuracy of DSHN. DSHN shows no signs of oversmoothing, as accuracy improves as we add layers. Performance also increases with a higher stalk dimension $d$, underscoring the additional expressive power associated to cellular sheaves. This stands in clear contrast to HGNN, whose accuracy steadily deteriorates with depth. This is in line with the observations in Bodnar et al. [2022] for graphs: leveraging our Directed Sheaf Hypergraph Laplacian, built with $d\times d$ restriction maps to transport features between nodes and hyperedges, enriches local variability rather than collapsing it. By projecting node features onto hyperedges (and back), the model retains discriminative power across neighborhoods.
@@ B.3 Extended results
In this subsection, we report the complete table of results for this work, which were not reported in the main just due to space limitations. As one can observed from Table 3, DSHN and DSHNLight consistently outperform the baselines taken from both the directed and undirected hypergraph learning literature on 10 out of 12 of the considered real-world datasets.

### v2

@@ A.1 Spectral Properties
By definition, we have $\mathbf{L}^{\vec{\mathcal{F}}}_{N}:=\mathbf{I}_{nd}-\mathbf{Q}^{\vec{\mathcal{F}}}_{N}$, with $\mathbf{Q}^{\vec{\mathcal{F}}}_{N}:=\mathbf{D}_{V}^{-\frac{1}{2}}{\mathbf{B}^{(q)}}^{\dagger}\mathbf{D}_{E}^{-1}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}$.
$\mathbf{Q}^{\vec{\mathcal{F}}}_{N}$ can be factored as
$\mathbf{Q}^{\vec{\mathcal{F}}}_{N}=\left(\mathbf{D}_{V}^{-\frac{1}{2}}{\mathbf{B}^{(q)}}^{\dagger}\mathbf{D}_{E}^{-\frac{1}{2}}\right)\left(\mathbf{D}_{E}^{-\frac{1}{2}}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}\right)=\left(\mathbf{D}_{E}^{-\frac{1}{2}}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}\right)^{\dagger}\left(\mathbf{D}_{E}^{-\frac{1}{2}}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}\right).$
@@
Since $\mathbf{L}^{\vec{\mathcal{F}}}_{N}:=\mathbf{I}_{nd}-\mathbf{Q}^{\vec{\mathcal{F}}}_{N}$, it follows that the spectrum of $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is upper-bounded by 1, which concludes the proof. ∎
@@ A.2 Generalization Properties
In the $2$-uniform case, every hyperedge $e$ contains exactly two nodes (i.e., $\delta_{e}=2$). Consider the general expression of the unnormalized Laplacian given in Eq. 3. Since the graph has no directions, $\mathcal{S}^{(0)}_{u\trianglelefteq e}=1$ for all $u\in V,e\in E$, and for any choice of the charge parameter $q$. As a result, the off-diagonal terms of $\mathbf{L}^{\vec{\mathcal{F}}}$ are real-valued (the diagonal ones always are).
In particular, when $\delta_{e}=2$ for all $e\in E$, $\mathbf{L}^{\vec{\mathcal{F}}}$ reads:
[…]
Thus, $\mathbf{L}^{\vec{\mathcal{F}}}$ precisely coincides with the Sheaf Laplacian of Hansen and Gebhart (2020) up to the multiplicative constant $\tfrac{1}{2}$.
When considering the case of a trivial Sheaf (i.e., when $\mathcal{F}_{v\trianglelefteq e}=1$), $\mathbf{L}^{\vec{\mathcal{F}}}$ coincides with the definition of the classical graph Laplacian $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $A$ is the adjacency matrix and $\mathbf{D}$ is the node degree matrix.
@@
The Magnetic Laplacian proposed by Zhang et al. (2021) is defined as
$\mathbf{L}^{(q)}\coloneqq\mathbf{D}_{s}-\mathbf{H}^{(q)}=\mathbf{D}_{s}-\mathbf{A}_{s}\odot\exp\!\bigl(i\,\mathbf{\Theta}^{(q)}\bigr),$
where $\mathbf{\Theta}^{(q)}$ denotes the phase matrix defined as
$\mathbf{\Theta}^{(q)}\coloneq 2\pi q\,(\mathbf{A}-\mathbf{A}^{\top})$
and $\mathbf{A}_{s}$ is the symmetrized adjacency matrix defined as
$\mathbf{A}_{s}:=\tfrac{1}{2}(\mathbf{A}+\mathbf{A}^{\top})$
and $\mathbf{D}_{s}$ is a diagonal matrix defined as
[…]
Entry-wise, $\mathbf{H}^{(q)}$ can be written as:
$\mathbf{H}^{(q)}_{uv}=\begin{cases}\tfrac{1}{2}e^{\,2\pi iq}&(u,v)\in E\\[3.0pt] \tfrac{1}{2}e^{-\,2\pi iq}&(v,u)\in E\\[3.0pt] 1&\{u,v\}\in E\\[3.0pt] 0&\text{otherwise}.\end{cases}$
In the directed, 2-uniform case, every hyperedge $e$ contains exactly two nodes ($\delta_{e}=2$). For every $e\in E$, the product $(\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e})^{\dagger}\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}$ can take one of the following three values:
@@
    $\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}=\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}=1\quad\Longrightarrow\quad\big(\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\big)^{\dagger}\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}=1.$
@@
    $\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}=e^{-2\pi iq},\qquad\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}=1\quad\Longrightarrow\quad\big(\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\big)^{\dagger}\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}=e^{+2\pi iq}.$
@@
    $\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}=1,\qquad\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}=e^{-2\pi iq}\quad\Longrightarrow\quad\big(\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\big)^{\dagger}\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}=e^{-2\pi iq}.$
@@
[…]
@@
$\mathbf{Q}^{\vec{\mathcal{F}}}={\mathbf{B}^{(q)}}^{\dagger}\,\mathbf{D}_{E}^{-1}\,\mathbf{B}^{(q)}\;=\;\mathbf{H}^{(q)},\text{ with }\mathbf{D}_{V}\;=\;\mathbf{D}_{s}.$
@@
$\textbf{L}^{\vec{\mathcal{F}}}=\mathbf{D}_{s}-\mathbf{H}^{(q)}=\mathbf{L}^{(q)}.$
Lastly, noticing that, by construction, the Sign-Magnetic Laplacian proposed in Fiorini et al. (2023) coincides with the Magnetic Laplacian when $q=\frac{1}{4}$, we conclude that our operator also generalizes the former. ∎
@@
In the unit-weight case, the Laplacian matrix proposed by Zhou et al. (2006) for undirected hypergraphs is defined as follows:
$\mathbf{\Delta}:=\mathbf{I}-\mathbf{Q}_{N}\qquad\text{ with }\mathbf{Q}_{N}:=\mathbf{D}_{V}^{-\frac{1}{2}}\mathbf{B}\mathbf{D}_{E}^{-1}\mathbf{B}^{\top}\mathbf{D}_{V}^{-\frac{1}{2}}.$
Since any undirected hypergraph be regarded as a special case of a directed hypergraph in which every hyperedge consists solely of tail nodes (or, equivalently, solely of head nodes), as shown in Eq. 4, in our proposed Laplacian matrix $\textbf{L}^{\vec{\mathcal{F}}}_{N}$ each product of two restriction maps reduces to a real weight of $1$, therefore contributing only to the real part of the operator. In particular, for a trivial sheaf where $\mathcal{F}_{v\trianglelefteq e}=1$, the incidence matrix $\mathbf{B}^{(q)}$ in Eq. 1 reduces to the transpose of binary incidence matrix $B$ of Zhou et al. (2006). ∎
@@
Our Normalized Directed Sheaf Hypergraph Laplacian $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$, component-wise reads:
$(\mathbf{L}^{\vec{\mathcal{F}}}_{N})_{uv}=\left\{\begin{array}[]{lr}\displaystyle\par\mathbf{I}_{d}-\mathbf{D}_{u}^{-1}\sum_{e:u\in e}\frac{1}{\delta_{e}}{\mathcal{F}}_{u\trianglelefteq e}^{\top}{\mathcal{F}}_{u\trianglelefteq e}&u=v\par\\ \displaystyle-\mathbf{D}_{u}^{-\frac{1}{2}}\big(\sum_{e:u,v\in e}\frac{1}{\delta_{e}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\big)\mathbf{D}_{v}^{-\frac{1}{2}}&u\neq v.\end{array}\right$.
@@
$(\mathbf{L}^{\vec{\mathcal{F}}}_{N})_{uv}=\left\{\begin{array}[]{ll}\displaystyle 1-\sum_{e:\,u\in e}\frac{1}{\mathbf{D}_{u}\,\delta_{e}},&u=v,\\[6.0pt] \displaystyle-\!\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u,v\in H(e)\\ \lor\;u,v\in T(e)\end{subarray}}\!\!\!\!\!\frac{1}{\delta_{e}}-i\left(\sum_{\begin{subarray}{c}e\in E\\ u\in T(e)\\ \land v\in H(e)\end{subarray}}\!\!\frac{1}{\delta_{e}}-\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u\in H(e)\\ \land v\in T(e)\end{subarray}}\!\!\frac{1}{\delta_{e}}\right)\frac{1}{\sqrt{\mathbf{D}_{u}}\sqrt{\mathbf{D}_{v}}},&u\neq v.\end{array}\right$. (14)
Such an expression coincides with the definition of the Generalized Directed Laplacian when considering $\mathbf{W}=\mathbf{I}$.∎
@@ B.1 Impact of charge parameter
The charge parameter $q$ controls how much each directed hyperedge contributes to the real and imaginary parts of the Directed Sheaf Hypergraph Laplacian. Larger values of $q$ place more directional information in the imaginary component, whereas smaller values reduce the directional contribution, emphasizing orientation-agnostic interactions in the real part. Because the dataset differ in how informative directionality is, the optimal $q$ is inherently data-dependent. In practice, a careful tuning of it is needed $q$ to select the value that yields the best performance, allowing either a partial or a full contribution of directional information to be encoded as needed. Table 2 reports the values chosen by our hyperparameter tuning procedure. As one can seen, for most datasets the hyperparameter tuning procedure sets a relatively high importance to directional information for each dataset, particularly for Telegram, Roman-empire and Synthetic datasets.
@@ B.2 Impact of stalk dimension and number of layers
As noted in Section 1, standard HGNNs are prone to oversmoothing: as network depth increases, node representations become indistinguishable and accuracy degrades. In Fig. 4, we study how depth and the stalk dimension $d$ affect the accuracy of DSHN. DSHN shows no signs of oversmoothing, as accuracy improves as we add layers. Performance also increases with a higher stalk dimension $d$, underscoring the additional expressive power associated to cellular sheaves. This stands in clear contrast to HGNN, whose accuracy steadily deteriorates with depth. This is in line with the observations in Bodnar et al. (2022) for graphs: leveraging our Directed Sheaf Hypergraph Laplacian, built with $d\times d$ restriction maps to transport features between nodes and hyperedges, enriches local variability rather than collapsing it. By projecting node features onto hyperedges (and back), the model retains discriminative power across neighborhoods.
@@ B.3 Extended Results
In this subsection, we report the complete table of results for this work, which were not reported in the main just due to space limitations. As can be observed from Table 3, DSHN, and DSHNLight consistently outperform the baselines taken from both the directed and undirected hypergraph learning literature on 10 out of 12 considered real-world datasets.